equation of state corresponding to 2gm h2
a)pv=rt    b)pv=rt/4  3)pv=rt/3   4)pv=rt/2
 

Grade:11

2 Answers

Harsh
6 years ago
The answer is (a)No of moles (n) = given mass/molar mass = 2/2=1molePV = NRTPut value of n in above equationPV = NRT
Musk
2 years ago
From the formula PV=nRT.   
From problem no of moles (n) = 2/2 =1 
Substitute 'n' in above equation we get 
 Answer as.       "  PV=RT "